40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2304     Entry time: Tue Mar 12 16:58:50 2019
Author: awade 
Type: Misc 
Category: WOPO 
Subject: Todo list W11 

These are things to get do this week on WOPO experiment:

☑️ Reinstall 50:50 fiber splitter into homodyne setup and go back to fiber launch of both ends of the HD directly onto photodetectors (rather than free space)

☐ Check visibility of HD by launching 1064 nm into both arms of HD using splitter and extra length on one arm, ramp laser frequency to get fringes and look at Vpp on each diode seperatly

☑️ Optimize subtraction of HD for max Common Mode Rejection (CMR) of LO amplitude noise.  Inject 3.21 kHz line into laser BNC port and minimize this peak on the subtracted output of the HD.

☐ Check 1064 nm -> 532 nm conversion in WOPO device to establish polarization basis for correct pol alignment into fiber (change 532 nm launch polarization if necessary), should be along the fast axis but for some reason this isn't in the datasheet explictly

☐ Double check how much power change there is from 2 pi modulation depth from PZT mounted 1064 nm mirror.  We don't want to over actuate on this element as it slightly misaignes into the fiber launch and causes some change in power, this is ok over a small range as the CMR of the homodyne will reject this.  We just want to be sure that this isn't the dominant effect that we are seeing at the output once we thing we should be seeing SQZ.


☑️Tick mark when done.

ELOG V3.1.3-